Use the model for the period of a pendulum, T, such that
T = 2π sqrt(L/g) where the length of the pendulum is L and the acceleration due to gravity is g.
If the gravity is 32 ft/s^2 and the period equals 9 s, find the length to the nearest in.
(12 in. = 1 ft).
Round your answer to the nearest in.

Answer by rothauserc(4718)

2 * pi * sqrt(L/g) = T
:
2 * (22/7) * sqrt(L/32) = 9
:
sqrt(L/32) = 9 * 7/44 = 1.4318
:
L/32 = 2.05
:
L = 65.6 feet
:
0.6 of 12 inches is 7.2 or approximately 7 inches
:
65 * 12 + 7 = 787 inches
: